1/2x+1/3y=13
1/5x=1/8y=5
solve the system of equations using the elimination of addition method.
Identify whether the system is intersecting, parallel or coincident lines.

Answer by ankor@dixie-net.com(22740)

1/2x + 1/3y = 13
1/5x + 1/8y = 5
multiply the 1st equation by 2, mult the 2nd equation by 5
x + 2/3y = 26
x + 5/8y = 25
---------------Subtraction eliminates x
2/3y - 5/8y = 1
mult equation by 24 to get rid of those annoying fractions
16y - 15y = 24
y = 24
Find x
x + 5/8(24) = 25
x + 15 = 25
x = 25 - 15
x = 10
:
Intersects at 10, 24
:
Check solutions in the 1st original equation
1/2x + 1/3y = 13
1/2(10) + 1/3(24) = 13
5 + 8 = 13, confirms our solution